P and Q each stand for whole numbers.
P+Q= 1000
P is 150 greater than Q.

Answer by fractalier(6550)

From
p + q = 1000 and
p = q + 150
we substitute the second equation into the first and get
q + 150 + q = 1000
2q + 150 = 1000
2q = 850
q = 425
p must then be 575.